# Enhancing Demand Prediction in Open Systems by Cartogram-aided Deep Learning

### v1

@@ I Introduction
However, predicting the temporal patterns of the rental and return is still challenging for the following reasons: (i) It is an open system—the total number of users fluctuates, and the installation and shutdown of the stations are more frequent than the conventional systems. (ii) There exists an imbalance of rental and return between stations. Due to such uncertainty, some machine learning models have been designed for the prediction of “rental” or “return” (collectively called “demand” in this study) at a station level  Lin et al. (2018); Yao et al. (2018); Zhang et al. (2020a). Specifically, to utilize temporal and spatial patterns, previous studies have adopted either of the recurrent neural network or convolution neural network (CNN) Zhai et al. (2021); Feng et al. (2020) or both simultaneously Wang et al. (2021, 2020). However, the CNN is limited to capturing information of adjacent regions only, naturally leading to dismissal of long-range correlations between geographically distant regions with similar characteristics such as floating population and facility density. This limitation can be solved by introducing a graph neural network (GNN) that uses graph information between regions, accompanying the demanding computation Kipf and Welling ; Lee et al. . To overcome computational complexity, many studies have adopted coarse-grained demand by aggregating some demands at stations for a given window Zhang et al. (2017); Li et al. (2019); Kim et al. (2019). Yet, the coarse-graining process could blur the regional characteristics (for instance, the distribution and pattern of demand).
@@
In this study, to secure both the lower complexity and higher accuracy, we introduce the cartogram approaches obtained by iterating the Voronoi tessellation Du et al. (1999). The cartogram is a distorted map based on a feature of interest. The map of stations’ locations is distorted by spreading the stations’ positions using the Voronoi tessellation, until the station density of each Voronoi cell, equivalent to the inverse size of the Voronoi polygon, becomes homogeneous. As a result, stations of similar characteristics get clustered nearby, which is evidenced by the correlation coefficient. Furthermore, the uniform spatial distribution enables us to predict the new demand of newly installed stations that did not appear in the training data, which has not been accomplished so far. The absence-in-training-data but presence-in-test-data has hindered the long-time scale prediction, yet our framework can infer the initial demand of a new station by the adjacent demands in the spirit of a mean-field approach.
@@
We explore the demand (rental and return) data of public bicycles in Seoul (the capital city of South Korea) sta : year 2018 for training data and 2019 for test data (prediction). To utilize and predict the spatio-temporal patterns, we employ a spatial-temporal convolutional graph attention (ST-CGA) network  Zhang et al. (2020b) that consists of mainly three parts as self-attention Vaswani et al. (2017), graph attention network Veličković et al. , and CNN. For the efficiency and improved performance, we consider three different time scales of an hour, a day, and a week, and modify the self-attention into the batch attention and the node-feature update in the graph attention network, compared with the ordinary ST-CGA network. In particular, the batch attention refers to various data at different times, which leads to contemplating the temporal correlation, and then we accomplish multiple prediction results at once with higher accuracy. We believe that our framework is applicable for predicting temporal patterns even for untrained spatial data.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce the empirical data and the ST-CGA model, utilizing the cartogram approaches. Especially, we focus on the modification in the ST-CGA model and the effects before and after applying to the cartogram idea. In Sec. III, we showcase the prediction performance overall and the initial demand prediction of a newly installed station. Lastly, we summarize this paper and provide a discussion in Sec. IV.
@@ II Data construction and Prediction method
In this study, we analyze and predict the spatio-temporal demand patterns of public bicycles in Seoul, Korea, which is definitely an open system. The rental or return patterns are spatially heterogeneous across the rental stations, as seen in Figs. 1(a) and 1(b), so we also use the CNN for prediction. However, the CNN only factors in the adjacent regions inherently limited by the size of its spatial window, which is called a filter, although the return-and-rental sometimes manifests itself between distant stations. Furthermore, some regions have similar characteristics, such as population density and land use, which could affect the demand pattern, even if the rental-and-return among them does not happen indeed. To contemplate such a long-range correlation, the graph neural network and relevant variants using network information have been developed. Among others, we exploit the ST-CGA network Zhang et al. (2020b) that embraces both characteristics of the convolutional and graph neural networks, which we also modify to enhance precision. In this section, we describe the conversion of the empirical data into a suitable form as input data and an entire architecture of the prediction model, focusing on our modification.
@@ II.1 Rental-and-return data in Seoul
Figure 1: The usage of public bicycles in Seoul. Snapshots of rentals (one of the two types of demand) (a) at 18:00, June 3rd, 2018 and (b) at 18:00, June 3rd, 2019. The sum $X$ of rentals in a cell on a grid is indicated by color in logarithmic scale. Note that a cell marked by a red arrow solely has a newly installed station in year 2019, of which the administrative-gu is Eunpyeong-gu. (c) A time series of rentals in year 2019. The mean value spanning a week is plotted as guidance to showcase the quasi-periodicity (we actually use the original sequence, not this mean value). As an illustrative example, we mark the $\tau_{d}$ points of rentals $X$ required to predict $X(t+\Delta t_{h})$, when considering a day resolution $\Delta t_{d}$.
We collect time series data for rentals and returns every hour from 0:00-0:59 on January 1, 2018 to 23:00-23:59 on December 31, 2019 sta . Thus, we have the time stamp [(365 days)$\times$(24 hours)] for two years, so $t\in\mathbb{T}$ with $\mathbb{T}=\{1,2,\cdots,8760\}$ for a given year (that is, $t=1$ stands for 0:00 on January 1, and $8760$ for 23:00 on December 31 of the same year). We will use the empirical data for the year 2018 as a training data set and those for the year 2019 as a test data set.
@@
To understand the characteristics of the temporal patterns, we display the averaged pattern of hourly rental, over cells and for year 2019 in Fig. 1(c). The quasi-periodic behaviors in different time scales are shown, e.g., rush hour/non-rush hour, weekday/weekend, and seasonal effect (although not shown here but straightforward to be expected), which is persuasive for taking the various time scales to seize this quasi-periodicity for training.
@@ II.2 Spatio-temporal convolutional graph attention network
Self-attention or batch attention: A set $\mathbb{X}^{r}(t)$ of demand vectors is used as input and all possible pairwise correlation are evaluated in a latent space to attain the predicted demand value $\hat{\mathbf{X}}(t+\Delta t_{h})\equiv[\hat{X}_{1}(t+\Delta t_{h}),\hat{X}_{2}(t+\Delta t_{h}),\cdots,\hat{X}_{MN}(t+\Delta t_{h})]^{T}$. It is called the self-attention; only referring to the current data itself, relevant to the pivot time $t$ in this study (i.e., $\mathbb{X}^{r}(t)$ tautologically). In the self-attention, the cell-to-cell (spatial) correlation and the temporal correlation within the time interval $\tau_{r}$ are concerned. The previous study Cheng et al. has revealed that the extended version of self-attention improves prediction performance. Referring to different data together is called batch attention. In this paper, the “different data” indicate the data at the different pivot time $t^{\prime}$ ($t\neq t^{\prime}$). In other words, adopting the batch attention means considering the mixture of the cell-to-cell and temporal correlation in a wide range of time stamps represented by different pivot times.
The number of the data to be referred corresponds to a batch size $B$, and we set this hyperparameter $B=16$. We randomly choose the $B$ different pivot times, and a bundle of $\mathbb{X}^{r}(t_{b})$’s (also $t_{b}\in\mathbb{T}$ and $b=0,1,\cdots,B-1$) becomes a new input unit. When $B=1$, it returns to the self-attention. Utilizing the batch attention enables us to attain the $B$ predicted values $\hat{\mathbf{X}}(t_{b}+\Delta t_{h})$ at once at the final stage. We adopt the batch attention but still use the terminology self-attention by convention.
This self-attention yields a $MNB\times L$ matrix $\mathbf{y}^{r}$ with $L$ being dimensions of the latent space and a hyperparameter (see Appendix A). The vector of the $u$-th row of the matrix is denoted as $\mathbf{y}^{r}_{u}$, where $u=i+b\times MN$, stores $L$ features of a cell $i$ at a pivot time $t_{b}$, which embodies all the influences of the other $v$-th rows, where $v=i^{\prime}+b^{\prime}\times MN$, calculated in the latent space. Although we cannot interpret the exact physical meaning of the features as always in the deep learning, this matrix is often called the correlation matrix since it carries some relatedness among cells across times. The detail is described in Appendix A.
@@
This modified update process has been verified in the previous study on an approximate personalized propagation of neural predictions Gasteiger et al. . We conjecture that the structure of graph attention network considered here may be similar to that of the approximate personalized propagation, supported by the improved performance.
Predicted value: The updated node features $\mathbf{y}_{i}^{\prime}$’s lead to the final output matrices $\mathbf{D}^{h}$, $\mathbf{D}^{d}$, or $\mathbf{D}^{w}$ of the CNN, whose important channels are determined by the feed forward neural network. Merging the three matrices, we accomplish a final output $\hat{\mathbf{X}}$ using another learnable importance matrix $\mathbf{W}$ computed as
@@
Measuring the mean squared loss in the learning process computed as
@@ II.3 Cartogram approaches to spread the stations
We pay attention to an open system confronted with a varying size of the system. The rental-and-return system also undergoes variation of demand either by the number of users or by the installation or shutdown of the stations. The absence of demand in a cell means null input data for training, which hinders us from predicting its future demand. To overcome this problem, we uniformly spread the stations across the map like a cartogram, so that empty cells do not exist. The cartogram is a thematic map that contains demographic or other features and yields the visually distorted geographic size, proportional to the relative level of feature of interest Sun and Li (2010); Gastner et al. (2005). In the sense that we visually distort a map suitably to our purpose – homogeneous distribution of the stations, the result of the following method using a tessellation is a cartogram based on the facility density.
We build up the process to spread the stations as follows, using a Voronoi cell that is a polygon containing only one point Aurenhammer (1991) as illustrated in Fig. 3: (i) Make the triangles fully packed in the plane by connecting the three nearest points (stations), without any cross lines and overlap among the triangles. (ii) For every triangle, form a circumcircle and then find the center of the circle. (iii) Form new polygons by connecting the centers of the circles without any cross lines and overlap among others, and position the points at the centers of the polygons that they belong to. We iterate the steps (i)-(iii) until there is no significant change in the spatial distribution of the stations, then any empty cell is absent in the new map divided into the $M\times N$ grid. The new cells after the sufficient iterations compose the input data $\mathbb{X}$ in Eq. (3). It can be seen that the areal distribution of the polygons becomes homogeneous after iteration in Figs. 3(c) and 3(d). The homogeneity allows us to predict demand of a newly installed station, in a similar spirit of a mean-field approach.
Figure 4: The similarity of the demand patterns of stations within a cell before [(a), (c)] and after [(b), (d)] using the cartogram (representatively, the rental for year 2019). (a), (b) Pearson’s correlation coefficient $\rho$ for every cell. (c), (d) The coefficient of variation $v$. The cells with $\rho=1$ and $\nu=0$ colored in gray have only one station each.
@@
with the mean $\mu_{i}(t)=\frac{1}{|\mathbb{R}_{i}|}\sum_{j\in\mathbb{R}_{i}}x_{j}(t)=\frac{1}{|\mathbb{R}_{i}|}X_{i}$ and standard deviation $\sigma_{i}^{2}(t)={\frac{1}{|\mathbb{R}_{i}|}\sum_{j\in\mathbb{R}_{i}}\left[x_{j}(t)-\mu(t)\right]^{2}}$ at time $t$ (||$\cdots$|| ⋯ | is a cardinality of a set). The coefficient of variance in the original map and the cartogram map for year 2019 as an example is displayed in Figs. 4(c) and 4(d). The mean $v$ is $1.0\pm 0.3$ for the original map, while $0.9\pm 0.2$ for the cartogram. Thus, our uniformization keeps the correlation and the dispersion of the demands similar to the original level, although the number of stations in cells and the number of cells are varied. We also remark that empty cells disappear on the map as seen in Figs. 4(b) and 4(d), which potentiates the demand prediction of the new station solely installed in the location marked by an arrow in Fig. 1(b).
@@ III Results
We train the empirical rental and return data separately for the year 2018 using the model framework in Sec. II, and then predict the demands for the year 2019 and compare the predicted values with the empirical data. We first analyze and the meaning of the attention score in Eq. (4), then verify the performance of this prediction framework and argue the demands of a station that had not existed in 2018 but have newly installed in 2019.
III.1 Analysis the attention score of the graph attention network
Figure 5: The analysis for the attention score between every cell pair in the graph attention network, for the training data (year 2018). (a) The yearly average attention score, or weighted adjacency matrix. The 255 cells correspond to a (coarse-grained) node, and the average attention score $\alpha_{ij}$ represents a weight of the link. The higher $\alpha_{ij}$ has a darker color. Clusters with low and high $\alpha$’s are surrounded by the two boxes in the upper left corner and lower right corner, respectively. The locations of the stations belonging to the nodes in the cluster with (b) the low attention scores and (c) the high attention scores are shown in the original map.
In the graph attention network, each cell that contains several stations is deemed a (coarse-grained) node, and the attention score $\alpha_{ij}$ in Eq. (4) stands for the link weight between nodes $i$ and $j$, which stores the long-range correlation between spatially distant nodes. The attention score matrix, or the weighted adjacency matrix, is shown in Fig. 5(a). The higher (lower) score $\alpha_{ij}$ represented as darker (lighter) red implies stronger (weaker) relatedness between two cells. We probe the two representative clusters of the lowest and highest weights encompassed by the purple and blue squares on the matrix, and mark the original positions of the stations in the map [Figs. 5(b) and 5(c)]. The stations belonging to the cluster with the lowest similarity are those that are almost distributed around the center. Otherwise, highly similar stations are scattered around the outskirts. It signifies that the outskirt regions do not have enough information to update node features by themselves. In other words, these regions need to get information from the other nodes (regions). That is why the outskirt regions have higher attention scores among them than the center regions of the city. Therefore, the network structure of our trained graph attention relates to the amount of information usage between nodes.

### v2

# Improving Demand Forecasting in Open Systems with Cartogram-Enhanced Deep Learning

@@ I Introduction
However, predicting the temporal patterns of the rental and return is still challenging for the following reasons: (i) It is an open system—the total number of users fluctuates, and the installation and shutdown of the stations are more frequent than the conventional systems. (ii) There exists an imbalance of rental and return between stations. Due to such uncertainty, some machine learning models have been designed for the prediction of “rental” or “return” (collectively called “demand” in this study) at a station level  Lin et al. (2018); Yao et al. (2018); Zhang et al. (2020a). Specifically, to utilize temporal and spatial patterns, previous studies have adopted either the recurrent neural network or convolution neural network (CNN) Zhai et al. (2021); Feng et al. (2020) or both simultaneously Wang et al. (2021, 2020). However, the CNN is limited to capturing information on adjacent regions only, naturally leading to the dismissal of long-range correlations between geographically distant regions with similar characteristics such as floating population and facility density. This limitation can be solved by introducing a graph neural network (GNN) that uses graph information between regions, accompanying the demanding computation Kipf and Welling ; Lee et al. . To overcome computational complexity, many studies have adopted coarse-grained demand by aggregating some demands at stations for a given window Zhang et al. (2017); Li et al. (2019); Kim et al. (2019). Yet, the coarse-graining process could blur the regional characteristics (for instance, the distribution and pattern of demand).
@@
In this study, to secure both the lower complexity and higher accuracy, we introduce the cartogram approaches obtained by iterating the Voronoi tessellation Du et al. (1999). The cartogram is a distorted map based on a feature of interest. The map of stations’ locations is distorted by spreading the stations’ positions using the Voronoi tessellation, until the station density of each Voronoi cell, equivalent to the inverse size of the Voronoi polygon, becomes homogeneous. As a result, stations of similar characteristics get clustered nearby, which is evidenced by the correlation coefficient. Furthermore, the uniform spatial distribution enables us to predict the new demand for newly installed stations that did not appear in the training data, which has not been accomplished so far. The absence-intraining-data but presence-in-test-data has hindered the long-time scale prediction. However, the successful prediction of brand-new demand naturally enables us to overcome such short-term predictions. The mean-field-like approach is in line with the fact that machine learning and physics have complementarily brought the advancement of each field Carleo et al. (2019).
@@
We explore the demand (rental and return) data of public bicycles in Seoul (the capital city of South Korea) sta : year 2018 for training data and 2019 for test data (prediction). To utilize and predict the spatio-temporal patterns, we employ a spatial-temporal convolutional graph attention (ST-CGA) network  Zhang et al. (2020b) that consists of mainly three parts as self-attention Vaswani et al. (2017), graph attention network Veličković et al. , and CNN. For efficiency and improved performance, we consider three different time scales of an hour, a day, and a week, and modify the self-attention into the batch attention and the node-feature update in the graph attention network, compared with the ordinary ST-CGA network. In particular, batch attention refers to various data at different times, which leads to contemplating the temporal correlation, and then we accomplish multiple prediction results at once with higher accuracy. We believe that our framework is applicable for predicting temporal patterns even for untrained spatial data.
@@
The rest of this paper is organized as follows: In Sec. II, we introduce the empirical data and the ST-CGA model, utilizing the cartogram approaches. Especially, we focus on the modification in the ST-CGA model and the effects before and after applying it to the cartogram idea. In Sec. III, we showcase the prediction performance overall and the initial demand prediction of a newly installed station. Lastly, we summarize this paper and provide a discussion in Sec. IV.
@@ II Data construction and Prediction method
In this study, we analyze and predict the spatio-temporal demand patterns of public bicycles in Seoul, Korea, which is definitely an open system. The rental or return patterns are spatially heterogeneous across the rental stations, as seen in Figs. 1(a) and 1(b), so we also use CNN for prediction. However, the CNN only factors in the adjacent regions inherently limited by the size of its spatial window, which is called a filter, although the return-and-rental sometimes manifests itself between distant stations. Furthermore, some regions have similar characteristics, such as population density and land use, which could affect the demand pattern, even if the rental-and-return among them does not happen indeed. To contemplate such a long-range correlation, the graph neural network and relevant variants using network information have been developed. Among others, we exploit the ST-CGA network Zhang et al. (2020b) that embraces both characteristics of the convolutional and graph neural networks, which we also modify to enhance precision. In this section, we describe the conversion of the empirical data into a suitable form as input data and an entire architecture of the prediction model, focusing on our modification.
@@ II.1 Rental-and-return data in Seoul
Figure 1: The usage of public bicycles in Seoul. Snapshots of rentals (one of the two types of demand) (a) at 18:00, June 3rd, 2018 and (b) at 18:00, June 3rd, 2019. The sum $X$ of rentals in a cell on a grid is indicated by color on a logarithmic scale. Note that a cell marked by a red arrow solely has a newly installed station in year 2019, of which the administrative-gu is Eunpyeong-gu. (c) A time series of rentals in the year 2019. The mean value spanning a week is plotted as guidance to showcase the quasi-periodicity (we actually use the original sequence, not this mean value). As an illustrative example, we mark the $\tau_{d}$ points of rentals $X$ required to predict $X(t+\Delta t_{h})$, when considering a day resolution $\Delta t_{d}$.
We collect time series data for rentals and returns every hour from 0:00-0:59 on January 1, 2018, to 23:00-23:59 on December 31, 2019 sta . The data is recorded at a minute-level unit, which exhibits noisy patterns due to factors such as human errors (record omissions) and drastic weather changes. To mitigate this noise, we aggregate the data from minute-level to hourly intervals. Thus, we have the time stamp [(365 days)$\times$(24 hours)] for two years, so $t\in\mathbb{T}$ with $\mathbb{T}=\{1,2,\cdots,8760\}$ for a given year (that is, $t=1$ stands for 0:00 on January 1, and $8760$ for 23:00 on December 31 of the same year). We will use the empirical data for the year 2018 as a training data set and those for the year 2019 as a test data set.
@@
To understand the characteristics of the temporal patterns, we display the averaged pattern of hourly rental, over cells and for the year 2019 in Fig. 1(c). The quasi-periodic behaviors in different time scales are shown, e.g., rush hour/non-rush hour, weekday/weekend, and seasonal effect (although not shown here but straightforward to be expected), which is persuasive for taking the various time scales to seize this quasi-periodicity for training.
@@ II.2 Spatio-temporal convolutional graph attention network
Self-attention or batch attention: A set $\mathbb{X}^{r}(t)$ of demand vectors is used as input and all possible pairwise correlation are evaluated in a latent space to attain the predicted demand value $\hat{\mathbf{X}}(t+\Delta t_{h})\equiv[\hat{X}_{1}(t+\Delta t_{h}),\hat{X}_{2}(t+\Delta t_{h}),\cdots,\hat{X}_{MN}(t+\Delta t_{h})]^{T}$. It is called the self-attention; only referring to the current data itself, relevant to the pivot time $t$ in this study (i.e., $\mathbb{X}^{r}(t)$ tautologically). In self-attention, the cell-to-cell (spatial) correlation and the temporal correlation within the time interval $\tau_{r}$ are concerned. The previous study Cheng et al. has revealed that the extended version of self-attention improves prediction performance. Referring to different data together is called batch attention. In this paper, the “different data” indicate the data at the different pivot time $t^{\prime}$ ($t\neq t^{\prime}$). In other words, adopting batch attention means considering the mixture of the cell-to-cell and temporal correlation in a wide range of time stamps represented by different pivot times.
The number of the data to be referred corresponds to a batch size $B$, and we set this hyperparameter $B=16$. We randomly choose the $B$ different pivot times, and a bundle of $\mathbb{X}^{r}(t_{b})$’s (also $t_{b}\in\mathbb{T}$ and $b=0,1,\cdots,B-1$) becomes a new input unit. When $B=1$, it returns to the self-attention. Utilizing batch attention enables us to attain the $B$ predicted values $\hat{\mathbf{X}}(t_{b}+\Delta t_{h})$ at once at the final stage. We adopt the batch attention but still use the terminology self-attention by convention.
This self-attention yields a $MNB\times L$ matrix $\mathbf{y}^{r}$ with $L$ being dimensions of the latent space and a hyperparameter (see Appendix A). The vector of the $u$-th row of the matrix is denoted as $\mathbf{y}^{r}_{u}$, where $u=i+b\times MN$, stores $L$ features of a cell $i$ at a pivot time $t_{b}$, which embodies all the influences of the other $v$-th rows, where $v=i^{\prime}+b^{\prime}\times MN$, calculated in the latent space. Although we cannot interpret the exact physical meaning of the features as always in deep learning, this matrix is often called the correlation matrix since it carries some relatedness among cells across times. The detail is described in Appendix A.
@@
This modified update process has been verified in the previous study on an approximate personalized propagation of neural predictions Gasteiger et al. . We conjecture that the structure of the graph attention network considered here may be similar to that of the approximate personalized propagation, supported by the improved performance.
Predicted value: The updated node features $\mathbf{y}_{i}^{\prime}$’s lead to the final output matrices $\mathbf{D}^{h}$, $\mathbf{D}^{d}$, or $\mathbf{D}^{w}$ of the CNN, whose important channels are determined by the feed-forward neural network. Merging the three matrices, we accomplish a final output $\hat{\mathbf{X}}$ using another learnable importance matrix $\mathbf{W}$ computed as
@@
Measuring the mean squared loss in the learning process is computed as
@@ II.3 Cartogram approaches to spread the stations
We pay attention to an open system confronted with a varying size of the system. The rental-and-return system also undergoes variations in demand either by the number of users or by the installation or shutdown of the stations. The absence of demand in a cell means null input data for training, which hinders us from predicting its future demand. To overcome this problem, we uniformly spread the stations across the map like a cartogram, so that empty cells do not exist. The cartogram is a thematic map that contains demographic or other features and yields the visually distorted geographic size, proportional to the relative level of feature of interest Sun and Li (2010); Gastner et al. (2005). In the sense that we visually distort a map suitably to our purpose – homogeneous distribution of the stations, the result of the following method using a tessellation is a cartogram based on the facility density.
We build up the process to spread the stations as follows, using a Voronoi cell that is a polygon containing only one point Aurenhammer (1991) as illustrated in Fig. 3: (i) Make the triangles fully packed in the plane by connecting the three nearest points (stations), without any cross lines and overlap among the triangles. (ii) For every triangle, form a circumcircle and then find the center of the circle. (iii) Form new polygons by connecting the centers of the circles without any cross lines and overlap among others and then find the center of the polygon. (iv) Move the positions of the points to the centers of the newly formed polygons that they belong to. We iterate the steps (i)-(iv) until there is no significant change in the spatial distribution of the stations, then any empty cell is absent in the new map divided into the $M\times N$ grid. The new cells after sufficient iterations compose the input data $\mathbb{X}$ in Eq. (3). It can be seen that the areal distribution of the polygons becomes homogeneous after iteration in Figs. 3(c) and 3(d). The homogeneity allows us to predict the demand of a newly installed station, in a similar spirit of a mean-field approach.
Figure 4: The similarity of the demand patterns of stations within a cell before [(a), (c)] and after [(b), (d)] using the cartogram (representatively, the rental for the year 2019). (a), (b) Pearson’s correlation coefficient $\rho$ for every cell. (c), (d) The coefficient of variation $v$. The cells with $\rho=1$ and $\nu=0$ colored in gray have only one station each.
@@
with the mean $\mu_{i}(t)=\frac{1}{|\mathbb{R}_{i}|}\sum_{j\in\mathbb{R}_{i}}x_{j}(t)=\frac{1}{|\mathbb{R}_{i}|}X_{i}$ and standard deviation $\sigma_{i}^{2}(t)={\frac{1}{|\mathbb{R}_{i}|}\sum_{j\in\mathbb{R}_{i}}\left[x_{j}(t)-\mu(t)\right]^{2}}$ at time $t$ (||$\cdots$|| ⋯ | is a cardinality of a set). The coefficient of variance in the original map and the cartogram map for the year 2019 as an example is displayed in Figs. 4(c) and 4(d). The mean $v$ is $1.0\pm 0.3$ for the original map, while $0.9\pm 0.2$ for the cartogram. Thus, our uniformization keeps the correlation and the dispersion of the demands similar to the original level, although the number of stations in cells and the number of cells are varied. We also remark that empty cells disappear on the map as seen in Figs. 4(b) and 4(d), which potentiates the demand prediction of the new station solely installed in the location marked by an arrow in Fig. 1(b). This significant increase in the input data point $X_{i}$, without violating the statistical level, is expected to enhance the accuracy of predictions.
@@ III Results
We train the empirical rental and return data separately for the year 2018 using the model framework in Sec. II, and then predict the demands for the year 2019 and compare the predicted values with the empirical data. We first analyze the meaning of the attention score in Eq. (4), then verify the performance of this prediction framework and argue the demands of a station that did not exist in 2018 but was newly installed in 2019.
III.1 Analysis of the attention score of the graph attention network
Figure 5: The analysis for the attention score between every cell pair in the graph attention network, for the training data (year 2018). (a) The yearly average attention score, or weighted adjacency matrix. The 255 cells correspond to a (coarse-grained) node, and the average attention score $\alpha_{ij}$ represents the weight of the link. The higher $\alpha_{ij}$ has a darker color. Clusters with low and high $\alpha$’s are surrounded by the two boxes in the upper left corner and lower right corner, respectively. The locations of the stations belonging to the nodes in the cluster with (b) the low attention scores and (c) the high attention scores are shown in the original map.
In the graph attention network, each cell that contains several stations is deemed a (coarse-grained) node, and the attention score $\alpha_{ij}$ in Eq. (4) stands for the link weight between nodes $i$ and $j$, which stores the long-range correlation between spatially distant nodes. The attention score matrix, or the weighted adjacency matrix, is shown in Fig. 5(a). The higher (lower) score $\alpha_{ij}$ represented as darker (lighter) red implies stronger (weaker) relatedness between two cells. We probe the two representative clusters of the lowest and highest weights encompassed by the purple and blue squares on the matrix, and mark the original positions of the stations in the map [Figs. 5(b) and 5(c)]. The stations belonging to the cluster with the lowest similarity are those that are almost distributed around the center. Otherwise, highly similar stations are scattered around the outskirts. It signifies that the outskirt regions do not have enough information to update node features by themselves. In other words, these regions need to get information from the other nodes (regions). That is why the outskirt regions have higher attention scores among them than the center regions of the city. Therefore, the network structure of our trained graph attention relates to the amount of information used between nodes.